Estimate the amount of Alcohol USP required to preserve the following solution:

100 ml of syrup containing  

                                                         10 g of sucrose

                                                         10 ml of propylene glycol
Answer #1

Given

Volume of preserved sucrose :10g (10g×0.53ml/g) = 5.3 ml

Volume occupied by sucrose is also 10g (10g×0.647 ml/g)=6.47 ml

Free water equivalent: 100ml_5.3_6.47=88.23ml

Volume of alcohol required to preserve the product is 88.23×18%=15.88

15.88 absolute alcohol ÷95%=16.71ml Alcohol USP or 17ml Alcohol USP